# Epicyclic motion of charged particles around a weakly magnetized Kiselev black hole

### v1

@@ 7 Conclusions
Once the effective potential is known, one can establish what types of orbits are possible. Contrary to Ernst spacetime, the potential (15) is vanishing on the horizons and it possesses off-equatorial critical points which correspond to saddle points, this feature being a result of quintessence. For a bound orbit to exist, the condition (23) should be satisfied and this leads to a range of the angular momentum given in (28) and to a maximum value of the radial coordinate of the turning point, defined in (29), which is depending only on the quintessence parameters $k$ and $w$. Also, in the figures 7 and 8, the key features are the shaded regions which correspond to the values of $L$ and $E$ which allow bound orbits.
A special attention is given to circular orbits. Since the potential vanishes on the horizons and it possesses at least one maximum value inbetween, it always exists un unstable circular orbit. On the other hand, stable circular orbits are possible for small values of $k$, lower than a critical value, and for the magnetic parameter in the range $b_{min}<b<b_{max}$, with $b_{min}$ and $b_{max}$ given in (45) and (46). The condition $b_{min}=b_{max}$ is satisfied for $r=r_{*}$, where $r_{*}$ is given in (29). The radial coordinate $r=r_{*}$ is particularly important since for $r$ above $r_{*}$ no stable orbits exist. The constrains on the quintessence parameter $k$ and the $r_{ISCO}$ ranges are collected in table 3. An important feature is that, due to quintessence, values of $r_{ISCO}$ bigger than the reference value $r=6M$ are possible. However, once the magnetic field comes into place, the ISCO radius shifts to lower values approaching the limiting value $r=3M$ (see the figure 6).
@@
In addition, we discuss the periastron and nodal precessions. The values and signs of $\Delta\phi$ and $\Delta\theta$ are depending not only on the magnetic parameter $b$, but also on the quintssence parameter $k$. As an interesting result, as $k$ increases, retrograde orbits may turn into prograde ones.
In the final subsection 6.3, we are giving some illustrative examples of trajectories for specific choices of the model’s parameters. In order to draw the particle’s trajectory by solving the equations of motion (13) using a numerical procedure, one must verify all the conditions previously discussed and fix suitable values for $B_{0}$, $k$ and $L$. The colored circles in the projections of the trajectories on the $xOy$ plane (see the figures 1821) are dividing physically important regions, as for example the green circle separates the region between curly and non-curly orbits, the gray circle corresponds to the change of sign of $\Delta\phi$, while the black circle corresponds to $r_{*}$ above which no stable trajectory exists. Similarly to Ernst spacetime, the particle’s trajectory may be curly or non-curly [8]. However, when the quintessence contribution is dominant, the unstable trajectory is curled outward the black hole and this behavior has no equivalence in Ernst spacetime where the trajectory is always curled toward the black hole as a result of the gravitational attraction.

### v2

@@ 7 Conclusions
From an astrophysical perspective, our investigation may serve as an effective framework for describing the motion of charged particles in realistic environments where supermassive or stellar-mass black holes surrounded by quintessence are also immersed in large-scale magnetic fields. Observations from type Ia supernovae [15] and microwave background radiation [39] indicate that the Universe expands at an accelerated rate, being driven by the mysterious dark energy characterized by a negative pressure. Numerous theoretical models have been proposed to explain the nature of the dark energy and one of the most prominent candidate is the quintessence. For realistic cosmological scenarios, the energy density of the quintessential fluid is expected to be very small [40, 41]. However, it alters the structure of the effective potential at large distances, modifies the stability conditions and influences particle trapping, escape channels and large-scale disk morphology. Observationally, the quantities most sensitive to the combined magnetic and quintessence effects are the ISCO radius and the associated periastron and nodal precessions. Moreover, the the radial and vertical epicyclic frequencies are directly connected to high-frequency quasi-periodic oscillations (HFQPOs) observed in X-ray binaries and to disk precession phenomena in accreting systems [42].
@@
Once the effective potential is known, one can establish what types of orbits are possible. Contrary to Ernst spacetime, the potential (18) is vanishing on the horizons and it possesses off-equatorial critical points which correspond to saddle points, this feature being a result of quintessence. For a bound orbit to exist, the condition (27) should be satisfied and this leads to a range of the angular momentum given in (32) and to a maximum value of the radial coordinate of the turning point, defined in (33), which is depending only on the quintessence parameters $k$ and $w$. Also, in the figures 9 and 10, the key features are the shaded regions which correspond to the values of $L$ and $E$ which allow bound orbits.
A special attention is given to circular orbits. Since the potential vanishes on the horizons and it possesses at least one maximum value inbetween, it always exists un unstable circular orbit. On the other hand, stable circular orbits are possible for small values of $k$, lower than a critical value, and for the magnetic parameter in the range $b_{min}<b<b_{max}$, with $b_{min}$ and $b_{max}$ given in (49) and (50). Outside this range whose width is decreasing as $k$ increases, no stable circular orbits exist. The condition $b_{min}=b_{max}$ is satisfied for $r=r_{*}$, where $r_{*}$ is given in (33). As we pointed out, the static radius $r=r_{*}$ is particularly important since it separates the region of stable orbits from the one where only unstable trajectories exist. The constrains on the quintessence parameter $k$ and the $r_{ISCO}$ ranges are collected in table 3. An important feature is that, due to quintessence, values of $r_{ISCO}$ bigger than the reference value $r=6M$ are possible. However, once the magnetic field comes into place, the ISCO radius shifts to lower values approaching the limiting value $r=3M$ (see the figure 8).
@@
In addition, we discuss the periastron and nodal precessions. The values and signs of $\Delta\phi$ and $\Delta\theta$ are depending not only on the magnetic parameter $b$, but also on the quintessence parameter $k$. As an interesting result, as $k$ increases, retrograde orbits may turn into prograde ones.
In the final subsection 6.3, we are giving some illustrative examples of trajectories for specific choices of the model’s parameters. In order to draw the particle’s trajectory by solving the equations of motion (16) using a numerical procedure, one must verify all the conditions previously discussed and fix suitable values for $B_{0}$, $k$ and $L$. The coloured circles in the projections of the trajectories on the $xOy$ plane (see the figures 2023) are dividing physically important regions, as for example the green circle separates the region between curly and non-curly orbits, the gray circle corresponds to the change of sign of $\Delta\phi$, while the black circle corresponds to $r_{*}$ above which no stable trajectory exists. Similarly to Ernst spacetime, the particle’s trajectory may be curly or non-curly [8]. However, when the quintessence contribution is dominant, the unstable trajectory is curled toward the black hole and this behavior has no equivalence in Ernst spacetime where the trajectory is always curled outward the black hole.
Acknowledgements
The authors would like to thank the anonymous Referee whose remarks and suggestions helped us to improve our manuscript.